OpenStudy (anonymous):

help please 1-cos^2θ/sinθ will give medal

OpenStudy (anonymous):

Do you know what sin^2 + cos^2 = ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

sin^2 + cos^2 = 1 Rewrite it with subtracting cos^2 from both sides

OpenStudy (anonymous):

1-cos

OpenStudy (anonymous):

1 - cos^2 = ?

OpenStudy (anonymous):

sin^2θ

OpenStudy (anonymous):

So your original equation was (1 - cos^2 theta)/sin theta what is your numerator = to?

OpenStudy (anonymous):

sin

OpenStudy (anonymous):

Sin^2 So it is sin^2/sin^2 which = ? I think you have the answer. I got to run.

OpenStudy (anonymous):

it is 1

OpenStudy (anonymous):

am i correct @Nnesha

OpenStudy (anonymous):

is it 1

Nnesha (nnesha):

there is sin the at the denominator right ?

OpenStudy (anonymous):

yes

Nnesha (nnesha):

\[\frac{ 1-\cos^2 \theta }{ \sin \theta }\]

Nnesha (nnesha):

as mentioned above replace 1-cos^2 tehta with sin^2 theta

OpenStudy (anonymous):

1-cos

Nnesha (nnesha):

what do you mean ?

Nnesha (nnesha):

it's 1 -cos^2theta

Nnesha (nnesha):

\[\frac{\color{Red}{ 1-\cos^2 \theta }}{ \sin \theta }\] according to the identity \[\rm \sin^2 \theta +\cos^2 \theta =1 \]\[\sin^2 \theta =1-\cos^2\] right

OpenStudy (anonymous):

yes

Nnesha (nnesha):

i think he made a typo at the end it should be \[\rm \frac{ \sin^2 \theta }{ \sin \theta }\]

Nnesha (nnesha):

now simplify that

OpenStudy (anonymous):

sinθ

Nnesha (nnesha):

right

OpenStudy (anonymous):

is that correct

OpenStudy (anonymous):

thanks

Nnesha (nnesha):

np

OpenStudy (anonymous):

@Nnesha @bluenile955 I did make a typo. Nnesha is correct.